40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4849     Entry time: Tue Jun 21 19:54:33 2011
Author: Suresh 
Type: Update 
Category: Green Locking 
Subject: LightWave NPRO power supply shifted to ETMY end table 

The Lightwave NPRO power supply which is being shared between the AS table and the ETMY table has been shifted back to the ETMY table

The current to the laser is set at 1.5A.  The laser output is 200mW at this current level.

ELOG V3.1.3-